OpenStudy (anonymous):

solve by completing the square: x^2+11x-1=0

OpenStudy (anonymous):

\[x^2+11x=1\] \[(x+\frac{11}{2})^2=1+\frac{121}{4}=\frac{125}{4}\] \[x+\frac{11}{2}=\pm\frac{\sqrt{125}}{2}\] \[x+\frac{11}{2}=\pm\frac{5\sqrt{5}}{2}\] \[x=-\frac{11\pm5\sqrt{5}}{2}\]

OpenStudy (anonymous):

sorry last step should be \[x=\frac{-11\pm5\sqrt{5}}{2}\] i had the minus sign in the wrong place

OpenStudy (anonymous):

x^2+11x-1=0 tric to complete square (1) divide by the coefficient of x^2 x^2+11x-1=0 (2) move the constant to the other side x^2+11x=1 (3) tahe half of the coefficient of x,square it and add it to both side x^2+11x+(11/2)^2=1+(11/2)^2 x^2+11x+(11/2)^2=125/4 x^2+2.11/2X+(11/2)^2=125/4 (x+11/2)^2=125/4 sqrting both side x+11/2=5sqrt5/2 therefore x=[-11(+-) *5 sqrt5]/2

OpenStudy (anonymous):